OpenStudy (anonymous):

zena spent 5 3/5 for 3 1/3 gallons of gas. how much did one gallon of gas cost?

OpenStudy (anonymous):

change them to improper fractions, then divide them to find the cost of a gallon

OpenStudy (anonymous):

i dont get it?

OpenStudy (anonymous):

5= 25/5 5 3/5 = 25/5+3/5

OpenStudy (anonymous):

what happend to the 1/3

OpenStudy (anonymous):

calculate it by yourself